Grade 12th passMechanics

an object is moving in a circle with constant speed v magnitude of the rate of change of momentum is

Anupam Das
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Momentum of an object (p) is equal to the mass of the object (m) times velocity of the object (v).So, p = mv …… (1)From equation (1) the rate of change of momentum of the object will be,dp/dt = d/dt (mv)= m dv/ (2) dt ……Since the object is moving in a circle at constant speed v (v = constant), therefore the term dv/dt in equation (2) will be equal to zero.So, dv/dt = 0 …… (3)Putting the value of equation (3) in equation (2), we get,dp/dt = m dv/dt= m×0= 0 …… (4)From equation (4) we observed that, the magnitude of the rate of change of momentum of the object is zero.